OpenStudy (anonymous):

Solve the equation using the Zero Product Property. -2x(5x-2)=0

OpenStudy (alexandervonhumboldt2):

x=0 x=2/5

OpenStudy (anonymous):

Let's solve your equation step-by-step. (-2x)(5x-2)=0 Step 1: Simplify both sides of the equation. -10x2+4x=0 Step 2: Factor left side of equation. 2x(-5x+2)=0 Step 3: Set factors equal to 0. 2x=0 or -5x+2=0 x=0 or x= 2 5